# Information-Consistent Language Model Recommendations through Group Relative Policy Optimization

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in business-critical domains such as finance, education, healthcare, and customer support, where users expect consistent and reliable recommendations. Yet LLMs often exhibit variability when prompts are phrased with minor differences, even when semantically equivalent. Such inconsistency undermines trust, complicates compliance, and disrupts user experience. While personalization is desirable in certain contexts, many enterprise scenarios—such as HR onboarding, customer support, or policy disclosure—require invariant information delivery regardless of phrasing or prior conversational history. Existing approaches, including retrieval-augmented generation (RAG) and temperature tuning, improve factuality or reduce stochasticity but cannot guarantee stability across equivalent prompts. In this paper, we propose a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) to directly optimize for consistency. Unlike prior applications of GRPO, which have been limited to reasoning and code generation, we adapt GRPO to enforce stability of information content across groups of semantically equivalent prompts. We introduce entropy-based helpfulness and stability rewards, treating prompt variants as groups and resetting conversational context to isolate phrasing effects. Experiments on investment and job recommendation tasks show that our GRPO-trained model reduces variability more effectively than fine-tuning or decoding-based baselines. To our knowledge, this is a novel application of GRPO for aligning LLMs toward information consistency, reframing variability not as an acceptable feature of generative diversity but as a correctable flaw in enterprise deployments.
@@ 1 INTRODUCTION
Solutions such as RAG lewis2020rag have been proposed as a way to support consistency by grounding answers in external knowledge. While effective in many contexts, RAG does not fully eliminate inherent inconsistencies in LLM behavior. For instance, consider two applicants preparing for the same job interview and querying an LLM-powered assistant: one asks, “What are the most common data structures I should review?” while another asks, “Which data structures should I prepare for in coding interviews?” Even when both queries are backed by the same retrieved documents, the LLM may generate divergent lists or emphasize different topics. Such inconsistencies create confusion and reduce trust in the system, especially in high-stakes contexts like interview preparation. Literature has often embraced this variability as an acceptable property of generative models, framing it as diversity in responses holtzman2019curious and proposing workarounds such as RAG or temperature tuning to control stochasticity. However, this acceptance does not resolve the fundamental issue. Simply lowering temperature or relying on retrieval does not guarantee that semantically equivalent prompts will produce consistent outputs. For many business and educational applications, overlooking this inconsistency is untenable.
@@
In addition to operational and reputational incentives, legal risk increasingly drives the demand for consistency in AI systems. Recent lawsuits highlight how inconsistent model behavior can expose organizations to liability. In Mobley v. Workday  hklaw2025workdaybias and Harper v. Sirius XM  fisherphillips2024sirius, allegations of hiring bias highlight how inconsistent or opaque decision patterns can be interpreted as unlawful disparate treatment. Cases involving chatbot misinformation, such as the British Columbia tribunal’s finding of liability for an AI-generated misstatement  americanbar2024chatbotliability, further illustrate that even occasional unreliability can create legal exposure. Likewise, the wrongful-death claim in Raine v. OpenAI  raine2024openai points to the risks posed by inconsistent adherence to safety-critical behaviors. These examples demonstrate that inconsistency undermines fairness, defensibility, and regulatory compliance, turning consistency from a technical preference into a legal imperative.
@@ 2 LITERATURE REVIEW
LLMs often exhibit sensitivity to superficial differences in prompts. Even when inputs are semantically equivalent, minor changes in form can cause divergent responses. Gan et al. gan2023promptrobustness document cross-lingual prompt sensitivity, while Sharma et al. sharma2024sensitivity propose explicit measures for quantifying consistency across paraphrases. Beyond phrasing effects, Liu et al. liu2024logicalconsistency highlight logical invariances—such as negation or transitivity—as proxies for stable reasoning, emphasizing that inconsistency is a pervasive issue for model reliability.
@@
A number of low-cost fixes attempt to mitigate inconsistency through decoding controls. Temperature tuning is frequently used to reduce variability, but work by Zhang et al. zhang2024temperature and Schmalbach schmalbach2025temp0 shows that deterministic decoding (e.g., setting temperature to zero) does not guarantee identical completions across semantically equivalent inputs. Thus, simple sampling adjustments cannot enforce true information consistency.
@@
RAG provides a popular strategy to stabilize outputs by grounding them in external sources lewis2020rag. While RAG improves factuality and reduces hallucination, it does not guarantee consistent outputs across paraphrased prompts. Even when retrieval results are held constant, divergence arises from stochasticity in generation and retrieval noise. This limitation is evident in contexts such as interview preparation, where two semantically equivalent queries may yield different recommended topics despite access to the same evidence. Consequently, RAG is complementary but insufficient for enforcing invariance.
@@
Fine-tuning has been explored as a more direct means of promoting consistency. Sun et al. sun2023zeroshotrobustness find that instruction-tuned models improve robustness overall but still exhibit notable brittleness under rephrasings. Zhao et al. zhao2024consistencyalignment propose a two-stage method combining instruction augmentation with consistency alignment training, showing reductions in paraphrase-induced variability. Wu et al. wu2024promptperturbation introduce prompt perturbation consistency learning, using divergence-based regularization to align outputs across perturbed variants. Raj et al. raj2025cog extend this with Chain of Guidance (CoG), generating synthetic variants for fine-tuning to enforce semantic agreement. These studies demonstrate that fine-tuning can help, but their approaches are primarily data-augmentation or loss-based, without making consistency the primary optimization goal. Recent approaches have addressed information consistency in retrieval-augmented generation (RAG) systems by optimizing similarity across paraphrased queries, often evaluated relative to retrieved evidence or task-specific ground truth answers hamman2025improving. While effective for knowledge-grounded QA, such formulations assume a well-defined notion of correctness. Our work instead focuses on business-critical recommendation and advisory settings, where ground truth is inherently ambiguous and the primary requirement is invariance of information content across user characteristics (e.g., gender or other attributes), independent of phrasing or retrieval context. This distinction motivates the need for consistency objectives that are independent of retrieval context and ground-truth supervision, particularly in enterprise-facing deployments.
@@
Despite these advances, important gaps remain. Most approaches either measure consistency post-hoc or encourage it indirectly. Few works define consistency as a direct training objective, and reinforcement learning has rarely been applied to explicitly minimize cross-variant dispersion. Furthermore, much of the literature evaluates on classification or reasoning tasks, leaving generative recommendation settings—such as investment or job advising—underexplored. Demographic attributes like gender have been used to highlight unequal opportunities salinas2023unequal, but less often as structured testbeds for prompt consistency itself.
@@ 2.1 Is Consistency Always Desirable in Business Scenarios?
For example, adaptive tutoring systems or AI-assisted learning environments benefit from adjusting responses to individual knowledge states, pacing preferences, or learning goals. A recent meta-analysis confirms that AI-driven personalized learning interventions yield consistently positive effects on student achievement across diverse contexts hu2024effect. Likewise, in healthcare, medical AI systems are expected to tailor guidance based on clinically relevant variables—such as age, comorbidities, or pregnancy status—where differentiation is justified by medical evidence rather than unintended demographic disparity who2024role_digitalhealth_women.
However, consistency becomes critical when the task involves factual correctness, safety, compliance, or fairness obligations. In such cases, equivalent inputs—regardless of how they are phrased or who provides them—should yield invariant outputs. Examples include banking disclosures, loan eligibility decisions, scientific or mathematical facts, and compliance-driven policies. This distinction aligns with principles in the NIST AI Risk Management Framework and the EU AI Act, which emphasize trustworthiness, reliability, and non-discrimination as essential for high-risk systems schwartz2022nist1270; euai2025regulation.
@@ 2.2 Balancing Fairness, Cultural Context, and Personalization
Ethical AI frameworks (e.g., UNESCO’s Recommendation on AI Ethics and the OECD AI Principles) advise that differentiation should be user-driven and contextually justified, not emergent from model inconsistencies unesco2022aiprinciples; oecd2019aiprinciples. This means cultural adaptation may be acceptable when explicitly requested by users (e.g., dietary guidance sensitive to religious practices) but not when identical prompts from users of different demographics elicit divergent factual or prescriptive answers.
@@ 2.3 The Boundary Between Personalization and Consistency
Our work addresses these gaps by leveraging Group Relative Policy Optimization (GRPO) 2024deepseekmath; guo2025deepseekr1; fan2025posteriorgrpo; robeyns2025grpocodequality. GRPO was originally introduced in the DeepSeekMath project as a lightweight alternative to PPO for reasoning tasks, where multiple samples per prompt are aggregated to stabilize mathematical or logical problem-solving 2024deepseekmath. Follow-up work has extended GRPO to scale reinforcement learning for general reasoning (DeepSeek-R1) guo2025deepseekr1, reward intermediate reasoning traces in code generation (Posterior-GRPO) fan2025posteriorgrpo, and improve code quality through reward shaping robeyns2025grpocodequality. To date, however, GRPO has been confined to reasoning and programming domains, focusing on factual accuracy and reasoning efficiency rather than enforcing stability across semantically equivalent prompts.
@@ 3.3 Prompt Variation and Consistency Testing (Gender Case Study)
Achieving consistency requires identifying and mitigating sources of unwanted variation in model outputs. One well-documented source of variation arises from the inclusion of demographic attributes in prompts. Even when irrelevant to the task, attributes such as gender or nationality can cause an LLM’s responses to diverge. Prior work has shown that language models can shift their outputs simply based on the presence of demographic cues, such as gender pronouns or nationality, resulting in a significant impact on the distribution of results salinas2023unequal.
@@
For our baseline experiment, we used the unsloth Llama-3 1B Instruct model meta2024llama3.2-1b-instruct; unsloth2024llama3.2-1b-instruct, which is instruction-tuned but not explicitly optimized for consistency. We posed prompts such as:
@@
and compared the resulting recommendations. A consistent model should recommend similar types of universities or courses regardless of the gender phrasing in the prompt. If systematic differences arise—such as favoring different fields or institutions between the two variants—it indicates inconsistency in the model’s outputs. This setup follows template-based testing strategies proposed in recent literature salinas2023unequal, where pronouns serve as controlled proxies for variation in prompt wording.
@@ 4.2 Reward Functions (Theory)
We operationalize the objective through two complementary rewards—Helpfulness (information richness) and Consistency (stability across variants)—combined into a single scalar objective optimized through Group Relative Policy Optimization (GRPO) 2024deepseekmath; guo2025deepseekr1; fan2025posteriorgrpo.
@@ 4.2.1 Helpfulness (Information) Reward
4.2.2 Consistency (Stability) Reward
@@
$F_{\text{norm}}=1-\frac{1}{K}\sum_{i<j}\frac{\left|H(r(q^{(i)}))-H(r(q^{(j)}))\right|}{\mathrm{MAX\_GAP}}.$
where
$\mathrm{MAX\_GAP}=\max_{1\leq k\leq n}\left|H_{k}-H_{k+n}\right|$
@@ 4.2.3 Composite Objective
where $\alpha$ controls emphasis on information richness and $\beta$ emphasizes stability. In high-stakes or fairness-sensitive domains, $\beta$ can be prioritized to make stability the dominant objective prabhune2025llms.
4.3 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) is a policy-gradient method designed for optimizing groups of samples per prompt 2024deepseekmath; guo2025deepseekr1; fan2025posteriorgrpo; robeyns2025grpocodequality. It extends the Proximal Policy Optimization (PPO) framework by computing advantage estimates relative to a group mean rather than an individual baseline, thereby aligning the optimization toward minimizing intra-group variance.
@@ 4.3 Group Relative Policy Optimization (GRPO)
Standard reinforcement learning (e.g., PPO or DPO) rewards single-sample performance and lacks an explicit term to minimize cross-sample variance. GRPO’s grouped formulation directly encodes variance minimization as part of the learning signal, making it particularly suitable for enforcing informational invariance across semantically equivalent inputs. This approach also generalizes beyond demographic groups to linguistic paraphrases, regional variants, or domain-specific phrasing differences fan2025posteriorgrpo; robeyns2025grpocodequality.
@@ 5.1 Dataset
We evaluated the proposed GRPO-based consistency framework on real-world investment and job recommendation prompts from the dataset shared by prabhune2025llms. This dataset contains semantically equivalent questions posed with different demographic markers. IT consists of 870 gendered questions derived from 400+ real user inquiries collected from public forums such as Reddit, Quora, and MarketWatch, spanning four business-relevant domains: Jobs, Education, Investment, and Health. Each real-world question was manually reviewed, filtered, and paired with semantically equivalent male and female variants to isolate the effect of the gender attribute. For this study, we focus on male/female prompt variants as a representative example of group-level differences.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly deployed in business-critical domains such as finance, education, healthcare, and customer support, where users expect consistent and reliable recommendations. Yet LLMs often exhibit variability when prompts are phrased with minor differences, even when semantically equivalent. Such inconsistency undermines trust, complicates compliance, and disrupts user experience. While personalization is desirable in certain contexts, many enterprise scenarios, such as HR onboarding, customer support, or policy disclosure, require invariant information delivery regardless of phrasing or prior conversational history. Existing approaches, including retrieval-augmented generation (RAG) and temperature tuning, improve factuality or reduce stochasticity, but cannot guarantee stability across equivalent prompts. In this paper, we propose a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) to directly optimize for consistency. Unlike prior applications of GRPO, which have been limited to reasoning and code generation, we adapt GRPO to enforce the stability of information content across groups of semantically equivalent prompts. We introduce entropy-based helpfulness and stability rewards, treating prompt variants as groups and resetting conversational context to isolate phrasing effects. Experiments on investment and job recommendation tasks show that our GRPO-fine-tuned model reduces variability compared to the baseline LLM model. To our knowledge, this is a novel application of GRPO for aligning LLMs toward information consistency, reframing variability not as an acceptable feature of generative diversity, but as a correctable flaw in enterprise deployments.
@@ 1 INTRODUCTION
Solutions such as RAG Lewis et al. (2020) have been proposed as a way to support consistency by grounding answers in external knowledge. While effective in many contexts, RAG does not fully eliminate inherent inconsistencies in LLM behavior. For instance, consider two applicants preparing for the same job interview and querying an LLM-powered assistant: one asks, “What are the most common data structures I should review?” while another asks, “Which data structures should I prepare for in coding interviews?” Even when both queries are backed by the same retrieved documents, the LLM may generate divergent lists or emphasize different topics. Such inconsistencies create confusion and reduce trust in the system, especially in high-stakes contexts like interview preparation. Literature has often embraced this variability as an acceptable property of generative models, framing it as diversity in responses Holtzman et al. (2019) and proposing workarounds such as RAG or temperature tuning to control stochasticity. However, this acceptance does not resolve the fundamental issue. Simply lowering temperature or relying on retrieval does not guarantee that semantically equivalent prompts will produce consistent outputs. For many business and educational applications, overlooking this inconsistency is untenable.
@@
In addition to operational and reputational incentives, legal risk increasingly drives the demand for consistency in AI systems. Recent lawsuits highlight how inconsistent model behavior can expose organizations to liability. In Mobley v. Workday  Holland & Knight LLP (2025) and Harper v. Sirius XM  Fisher Phillips LLP (2024), allegations of hiring bias highlight how inconsistent or opaque decision patterns can be interpreted as unlawful disparate treatment. Cases involving chatbot misinformation, such as the British Columbia tribunal’s finding of liability for an AI-generated misstatement  American Bar Association (2024), further illustrate that even occasional unreliability can create legal exposure. Likewise, the wrongful-death claim in Raine v. OpenAI  Raine and others (2024) points to the risks posed by inconsistent adherence to safety-critical behaviors. These examples demonstrate that inconsistency undermines fairness, defensibility, and regulatory compliance, turning consistency from a technical preference into a legal imperative.
@@ 2 LITERATURE REVIEW
LLMs often exhibit sensitivity to superficial differences in prompts. Even when inputs are semantically equivalent, minor changes in form can cause divergent responses. Gan et al. Gan and others (2023) document cross-lingual prompt sensitivity, while Sharma et al. Sharma and others (2024) propose explicit measures for quantifying consistency across paraphrases. Beyond phrasing effects, Liu et al. Liu and others (2024) highlight logical invariances—such as negation or transitivity—as proxies for stable reasoning, emphasizing that inconsistency is a pervasive issue for model reliability.
@@
A number of low-cost fixes attempt to mitigate inconsistency through decoding controls. Temperature tuning is frequently used to reduce variability, but work by Zhang et al. Zhang and others (2024) and Schmalbach Schmalbach (2025) shows that deterministic decoding (e.g., setting temperature to zero) does not guarantee identical completions across semantically equivalent inputs. Thus, simple sampling adjustments cannot enforce true information consistency.
@@
RAG provides a popular strategy to stabilize outputs by grounding them in external sources Lewis et al. (2020). While RAG improves factuality and reduces hallucination, it does not guarantee consistent outputs across paraphrased prompts. Even when retrieval results are held constant, divergence arises from stochasticity in generation and retrieval noise. This limitation is evident in contexts such as interview preparation, where two semantically equivalent queries may yield different recommended topics despite access to the same evidence. Consequently, RAG is complementary but insufficient for enforcing invariance.
@@
Fine-tuning has been explored as a more direct means of promoting consistency. Sun et al. Sun et al. (2024) find that instruction-tuned models improve robustness overall but still exhibit notable brittleness under rephrasings. Zhao et al. Zhao and others (2024) propose a two-stage method combining instruction augmentation with consistency alignment training, showing reductions in paraphrase-induced variability. Wu et al. Wu and others (2024) introduce prompt perturbation consistency learning, using divergence-based regularization to align outputs across perturbed variants. Raj et al. Raj and others (2025) extend this with Chain of Guidance (CoG), generating synthetic variants for fine-tuning to enforce semantic agreement. These studies demonstrate that fine-tuning can help, but their approaches are primarily data-augmentation or loss-based, without making consistency the primary optimization goal. Recent approaches have addressed information consistency in retrieval-augmented generation (RAG) systems by optimizing similarity across paraphrased queries, often evaluated relative to retrieved evidence or task-specific ground truth answers Hamman et al. (2025). While effective for knowledge-grounded QA, such formulations assume a well-defined notion of correctness. Our work instead focuses on business-critical recommendation and advisory settings, where ground truth is inherently ambiguous and the primary requirement is invariance of information content across user characteristics (e.g., gender or other attributes), independent of phrasing or retrieval context. This distinction motivates the need for consistency objectives that are independent of retrieval context and ground-truth supervision, particularly in enterprise-facing deployments.
@@
Despite these advances, important gaps remain. Most approaches either measure consistency post-hoc or encourage it indirectly. Few works define consistency as a direct training objective, and reinforcement learning has rarely been applied to explicitly minimize cross-variant dispersion. Furthermore, much of the literature evaluates on classification or reasoning tasks, leaving generative recommendation settings—such as investment or job advising—underexplored. Demographic attributes like gender have been used to highlight unequal opportunities Salinas et al. (2023), but less often as structured testbeds for prompt consistency itself.
@@ 2.1 Is Consistency Always Desirable in Business Scenarios?
For example, adaptive tutoring systems or AI-assisted learning environments benefit from adjusting responses to individual knowledge states, pacing preferences, or learning goals. A recent meta-analysis confirms that AI-driven personalized learning interventions yield consistently positive effects on student achievement across diverse contexts Hu (2024). Likewise, in healthcare, medical AI systems are expected to tailor guidance based on clinically relevant variables—such as age, comorbidities, or pregnancy status—where differentiation is justified by medical evidence rather than unintended demographic disparity World Health Organization Regional Office for Europe (2024).
However, consistency becomes critical when the task involves factual correctness, safety, compliance, or fairness obligations. In such cases, equivalent inputs—regardless of how they are phrased or who provides them—should yield invariant outputs. Examples include banking disclosures, loan eligibility decisions, scientific or mathematical facts, and compliance-driven policies. This distinction aligns with principles in the NIST AI Risk Management Framework and the EU AI Act, which emphasize trustworthiness, reliability, and non-discrimination as essential for high-risk systems Schwartz et al. (2022); European Parliament and Council of the European Union (2025).
@@ 2.2 Balancing Fairness, Cultural Context, and Personalization
Ethical AI frameworks (e.g., UNESCO’s Recommendation on AI Ethics and the OECD AI Principles) advise that differentiation should be user-driven and contextually justified, not emergent from model inconsistencies United Nations Educational, Scientific and Cultural Organization (2022); Organisation for Economic Co-operation and Development (2019). This means cultural adaptation may be acceptable when explicitly requested by users (e.g., dietary guidance sensitive to religious practices) but not when identical prompts from users of different demographics elicit divergent factual or prescriptive answers.
@@ 2.3 The Boundary Between Personalization and Consistency
Our work addresses these gaps by leveraging Group Relative Policy Optimization (GRPO) Shao et al. (2024); Guo and others (2025); Fan et al. (2025); Robeyns and Aitchison (2025). GRPO was originally introduced in the DeepSeekMath project as a lightweight alternative to PPO for reasoning tasks, where multiple samples per prompt are aggregated to stabilize mathematical or logical problem-solving Shao et al. (2024). Follow-up work has extended GRPO to scale reinforcement learning for general reasoning (DeepSeek-R1) Guo and others (2025), reward intermediate reasoning traces in code generation (Posterior-GRPO) Fan et al. (2025), and improve code quality through reward shaping Robeyns and Aitchison (2025). To date, however, GRPO has been confined to reasoning and programming domains, focusing on factual accuracy and reasoning efficiency rather than enforcing stability across semantically equivalent prompts.
@@ 3.3 Prompt Variation and Consistency Testing (Gender Case Study)
Achieving consistency requires identifying and mitigating sources of unwanted variation in model outputs. One well-documented source of variation arises from the inclusion of demographic attributes in prompts. Even when irrelevant to the task, attributes such as gender or nationality can cause an LLM’s responses to diverge. Prior work has shown that language models can shift their outputs simply based on the presence of demographic cues, such as gender pronouns or nationality, resulting in a significant impact on the distribution of results Salinas et al. (2023).
@@
For our baseline experiment, we used the unsloth Llama-3 1B Instruct model AI (2024a, b), which is instruction-tuned but not explicitly optimized for consistency. We posed prompts such as:
@@
and compared the resulting recommendations. A consistent model should recommend similar types of universities or courses regardless of the gender phrasing in the prompt. If systematic differences arise—such as favoring different fields or institutions between the two variants—it indicates inconsistency in the model’s outputs. This setup follows template-based testing strategies proposed in recent literature Salinas et al. (2023), where pronouns serve as controlled proxies for variation in prompt wording.
@@ 4.2 Reward Functions (Theory)
We operationalize the objective through two complementary rewards—Helpfulness (information richness) and Consistency (stability across variants)—combined into a single scalar objective optimized through Group Relative Policy Optimization (GRPO) Shao et al. (2024); Guo and others (2025); Fan et al. (2025).
@@ 4.2.1 Helpfulness (Information) Reward
Consistency (Stability) Reward
@@
$F_{\text{norm}}=1-\frac{1}{K}\sum_{k=1..K}\frac{\left|H(r(q_{k}^{(i)}))-H(r(q_{k}^{(j)}))\right|}{\mathrm{MAX\_GAP}}.$
where MAX_GAPMAX\_GAP is the max difference between the entropies of the two groups.
@@
where $\alpha$ controls emphasis on information richness and $\beta$ emphasizes stability. In high-stakes or fairness-sensitive domains, $\beta$ can be prioritized to make stability the dominant objective Prabhune et al. (2025).
Group Relative Policy Optimization (GRPO)
GRPO is a policy-gradient method designed for optimizing groups of samples per prompt Shao et al. (2024); Guo and others (2025); Fan et al. (2025); Robeyns and Aitchison (2025). It extends the Proximal Policy Optimization (PPO) framework by computing advantage estimates relative to a group mean rather than an individual baseline, thereby aligning the optimization toward minimizing intra-group variance.
@@
Standard reinforcement learning (e.g., PPO or DPO) rewards single-sample performance and lacks an explicit term to minimize cross-sample variance. GRPO’s grouped formulation directly encodes variance minimization as part of the learning signal, making it particularly suitable for enforcing informational invariance across semantically equivalent inputs. This approach also generalizes beyond demographic groups to linguistic paraphrases, regional variants, or domain-specific phrasing differences Fan et al. (2025); Robeyns and Aitchison (2025).
@@ 5.1 Dataset
We evaluated the proposed GRPO-based consistency framework on real-world investment and job recommendation prompts from the dataset shared by Prabhune et al. (2025). This dataset contains semantically equivalent questions posed with different demographic markers. IT consists of 870 gendered questions derived from 400+ real user inquiries collected from public forums such as Reddit, Quora, and MarketWatch, spanning four business-relevant domains: Jobs, Education, Investment, and Health. Each real-world question was manually reviewed, filtered, and paired with semantically equivalent male and female variants to isolate the effect of the gender attribute. For this study, we focus on male/female prompt variants as a representative example of group-level differences.